OpenStudy (anonymous):

solve for d 5d+2(2-d)=3(1+d)+1

OpenStudy (texaschic101):

5d + 2(2 - d) = 3(1 + d) + 1 -- distribute through the parenthesis 5d + 4 - 2d = 3 + 3d + 1 -- combine like terms 3d + 4 = 4 + 3d There is INFINITE SOLUTIONS for this